OpenStudy (anonymous):

if ln5=a and ln3=b write ln0.12 as an expression in a and b

OpenStudy (anonymous):

.12=2*3/2*5

OpenStudy (sgadi):

0.12=12/10=6/5=2*3/5 ln(2*3/5)=ln2+ln3-ln5=ln2+b-a

OpenStudy (anonymous):

sorry that is wrong. \[.12=\frac{2\times 3}{(2\times 5)^2}\]

OpenStudy (anonymous):

actually .12=12/100 but that is the right idea

OpenStudy (anonymous):

i mean sgadi has the right method

OpenStudy (anonymous):

so ln2+b-a is the answerÉ

OpenStudy (anonymous):

not what i get

OpenStudy (anonymous):

what you get im lost

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

\[.12=\frac{12}{100}\]

OpenStudy (anonymous):

now reduce this to \[\frac{12}{100}=\frac{3}{25}\]

OpenStudy (anonymous):

now it is easy

OpenStudy (anonymous):

ish

OpenStudy (anonymous):

but where does a and b come in

OpenStudy (anonymous):

\[ln(\frac{3}{25})=ln(\frac{3}{5^2})\]

OpenStudy (anonymous):

not there yet is is coming

OpenStudy (anonymous):

\[ln(\frac{3}{5^2})=ln(3)-ln(5^2)\]

OpenStudy (anonymous):

=\[ln(3)-2ln(5)\]

myininaya (myininaya):

ln(12/100)=ln(3/25)=ln3-ln5^2=ln3-2ln5=b-2a

OpenStudy (anonymous):

and there is your a and b. \[ln(3)-2ln(5)=a-2b\]

OpenStudy (anonymous):

the gimmick was to rewrite these in terms of ln(3) and ln(5)

myininaya (myininaya):

b=ln3 a=ln5 so its b-2a

OpenStudy (anonymous):

thanks you!